OpenStudy (anonymous):

what is (e^x)^2 equal

OpenStudy (anonymous):

\[\left( e ^{x} \right)^{2}\]

OpenStudy (amistre64):

a^2 = a*a (e^x)^2=e^x * e^x = e^(x+x)

OpenStudy (anonymous):

amistre64 you still their

OpenStudy (amistre64):

my internet connection is iffy

OpenStudy (anonymous):

you cant integrate e^(x^2) and the book has an answer its not like 2e^2x

OpenStudy (amistre64):

its not that you cant integrate it, you simply have to go about it in another manner using its taylor series equivalent

OpenStudy (amistre64):

there is no elementary function that we know of that wil derive down to e^(x^2)

OpenStudy (amistre64):

but if we construct a polynomial that is equivalent to it; then we can integrate the polynomial

OpenStudy (amistre64):

a taylor series is built by making a polynomial function that bends and moves in the same manner as e^(x^2)

OpenStudy (anonymous):

so i think your saying i can just use something like 2x^2+1

OpenStudy (amistre64):

something like that, but lets try to build the function :) f(x) = e^(x^2) f' = 2x e^(x^2) f''= e^(x^2) (4x^2+2) f''' = e^(x^2) (6x^4 ....) its tricky trying to do it on a keyboard

OpenStudy (amistre64):

but once youve got the derivatives that you want you see a pattern devlop that can be built in summation notations

OpenStudy (amistre64):

i think you can do a e^u and then sub in x^2 for u maybe

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=e%5E%28x%5E2%29 this might help :)